i need help to understand how to solve an equation with variables on both sides of the equal sign the problem is -v+5+6v=1+5v+3
Answer by Alan3354(69443)

-v+5+6v=1+5v+3
-------------------
Add or subtract until they're on one side.
-v+5+6v=1+5v+3
5v + 5 = 5v + 4
5 = 4
Not a good example.
There's no solution to this one.